OpenStudy (anonymous):

Your office is 60 km from your home .you drive the first 30 km at a constant 20 m/sec and the next 30 km at a constant 40.0 m/sec. what is your average speed for the entire speed for entire journey

OpenStudy (anonymous):

For the first part of the Journey the time taken t1= distance/speed t1 = 3000/20 = 150seconds For the second part of the journey t2 = 3000/40 = 75 seconds The average speed can be calculated by Ratio of total distance travelled and total time taken average speed = 6000/(t1 + t2) = 6000/(150+75) = 6000/225 = 26.666 meters per second

OpenStudy (anonymous):

why would 30km at 20m/sec be 3000/40=75 shouldnt it be 30000/40?